# Near-optimal discrimination of displaced squeezed binary signals using displacement, inverse-squeezing, and photon-number-resolving detection

### 17 Mar 2026 v2

@@ V Conclusion
Appendix A The Wigner Function of Displaced Squeezed Vacuum States
The displaced squeezed vacuum states discussed in this paper belong to single-mode Gaussian states, whose Wigner function is a bivariate Gaussian distribution fully characterized by two parameters: the displacement vector $\vec{d}$ and the covariance matrix $\mathbf{V}$. Given a displaced squeezed vacuum states denoted as $\left|\psi\right\rangle$, its quantum mechanical expression is
$\left|\psi\right\rangle=\left|\alpha,r\right\rangle=D(\alpha)S(r)\left|0\right\rangle,\qquad\alpha,r\in\mathbb{R}_{+}.$ (60)
The first parameter, the displacement vector $\vec{d}$, is a two-dimensional vector representing the translation in phase space relative to the origin. Its components correspond to the expectation values of the quadrature operators $\hat{X},\hat{P}$:
$\vec{d}=\begin{pmatrix}d_{x}\\ d_{p}\end{pmatrix}=\begin{pmatrix}\left\langle\hat{X}\right\rangle\\[2.0pt] \left\langle\hat{P}\right\rangle\end{pmatrix}$. (61)
For $\left|\psi\right\rangle$, $d_{x}=\sqrt{2}\,\mathrm{Re}(\alpha)=\sqrt{2}\alpha$ and $d_{p}=\sqrt{2}\,\mathrm{Im}(\alpha)=0$. The second parameter, the covariance matrix $\mathbf{V}$, is a $2\times 2$ real symmetric positive-definite matrix describing the “shape” and “size” of the quantum state in phase space. Its elements are variances and covariances:
$\mathbf{V}=\begin{bmatrix}V_{xx}&V_{xp}\\ V_{px}&V_{pp}\end{bmatrix}\\ =\begin{bmatrix}\left\langle(\Delta\hat{X})^{2}\right\rangle&\frac{1}{2}\left\langle\left\{\Delta\hat{X},\Delta\hat{P}\right\}\right\rangle\\[4.0pt] \frac{1}{2}\left\langle\left\{\Delta\hat{X},\Delta\hat{P}\right\}\right\rangle&\left\langle(\Delta\hat{P})^{2}\right\rangle\end{bmatrix}$, (62)
where $\Delta\hat{O}=\hat{O}-\left\langle\hat{O}\right\rangle$, and the anticommutator is defined as $\left\{\hat{A},\hat{B}\right\}=\hat{A}\hat{B}+\hat{B}\hat{A}$. For $\left|\psi\right\rangle$, the covariance matrix is
$\mathbf{V}=\begin{bmatrix}\frac{1}{2}e^{-2r}&0\\[2.0pt] 0&\frac{1}{2}e^{2r}\end{bmatrix}$, (63)
indicating that in phase space, the state is an ellipse (coherent states and vacuum states appear as circles), with fluctuations squeezed along the $X$-direction and expanded along the $P$-direction, while the overall phase-space area remains unchanged ($\det\mathbf{V}=\tfrac{1}{2}e^{-2r}\cdot\tfrac{1}{2}e^{2r}=\tfrac{1}{4}$).
Substituting $\vec{d}$ and $\mathbf{V}$ into the general formula yields the Wigner function for $\left|\psi\right\rangle$:
$\begin{aligned} &W(\vec{r}|\psi)=\frac{1}{2\pi\sqrt{\det\mathbf{V}}}\,\exp\left\{-\frac{1}{2}(\vec{r}-\vec{d})^{T}\mathbf{V}^{-1}(\vec{r}-\vec{d})\right\}\\ &\to W(x,p|\psi)=\frac{1}{\pi}\exp\left\{-\,\mathrm{e}^{2r}\big(x-\sqrt{2}\alpha\big)^{2}-\,\mathrm{e}^{-2r}p^{2}\right\},\end{aligned}$ (64)
where $\vec{r}=\begin{pmatrix}x\\ p\end{pmatrix}$ is a point in phase space, and $\det\mathbf{V}$ denotes the determinant of $\mathbf{V}$.
Appendix B The Helstrom bound and SQL of S-BPSK under phase diffusion condition
Phase diffusion noise is a communication channel noise, therefore theoretically both the Helstrom bound and the SQL for S-BPSK $\rho_{i}^{\text{pd}}$ (Eq. (33)) are affected. Specifically, the SQL expression in Eq. (4) should be rewritten as:
$P_{\text{SQL}}^{\text{S-BPSK}}\left(N,\sigma\right)$ $=\int_{0}^{+\infty}\mathrm{d}x\cdot P\left(x|\rho_{i}^{\text{pd}}\right)$ (65)
$=\int_{0}^{+\infty}\mathrm{d}x\int_{\mathbf{R}}\mathrm{d}\phi\cdot g_{\sigma}(\phi)\cdot P\left(x|\rho_{i,\phi}\right)$
$=\int_{0}^{+\infty}\mathrm{d}x\int_{\mathbf{R}}\mathrm{d}\phi\cdot g_{\sigma}(\phi)\int_{\mathbf{R}}\mathrm{d}p\cdot W\left(x,p|\rho_{i,\phi}\right)$
$=\int_{\mathbf{R}}\mathrm{d}\phi\cdot g_{\sigma}(\phi)\cdot P_{\text{SQL}}^{\text{S-BPSK}}(\phi)$
Here,
$P_{\text{SQL}}^{\text{S-BPSK}}(\phi)=\int_{0}^{+\infty}\mathrm{d}x\cdot P\left(x|\rho_{i,\phi}\right)=\int_{0}^{+\infty}\mathrm{d}x\int_{\mathbf{R}}\mathrm{d}p\cdot W\left(x,p|\rho_{i,\phi}\right)$ (66)
represents the SQL for the binary quantum states $\left\{\rho_{0,\phi},\rho_{1,\phi}\right\}$ (see Eq. (34)). $P(x|\rho_{i,\phi})$ is the homodyne probability of obtaining outcome $x$ given the state $\rho_{i,\phi}$. Note: The rotation operator $R_{\phi}$ is a Gaussian operator, thus $\rho_{i,\phi}$ remains a Gaussian state. Its Wigner function can be computed using the method described in Appendix A.
For the Helstrom limit of two states $\rho_{i}^{\text{pd}}$, Eq. (2) (applicable only to pure states) is no longer suitable and should be replaced with its equivalent form:
$P_{\text{HB}}^{\text{S-BPSK}}\left(N,\sigma\right)=\frac{1}{2}\left(1-\left\|P_{0}\rho_{0}^{\text{pd}}-P_{1}\rho_{1}^{\text{pd}}\right\|_{1}\right)$ (67)
where $\left\|A\right\|_{1}=\mathbf{Tr}\sqrt{A^{\dagger}A}$ denotes the trace norm, and $P_{0},P_{1}$ denote prior probabilities. This expression applies to the calculation of the Helstrom bound for binary states, whether pure or mixed.
Acknowledgements.
@@
Data Availability
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.
@@
No potential conflict of interest was reported by the authors

### 12 May 2026 v3

@@ Discussion
Methods
Threshold determination in non-ideal scenarios
For the phase-diffusion, receiver-thermal-noise, and transmission-loss analyses, a general closed-form proof of the monotonicity of the likelihood ratio $L(n)$ is not available. We therefore computed the conditional counting statistics $P(n|i)$ numerically and evaluated the error probability for all admissible threshold values within the detector-resolution window, $n_{\mathrm{th}}\in[1,M]$. The threshold was then chosen to minimize the error probability:
$P_{\mathrm{err,th}}^{\mathrm{ISK}}=\min_{n_{\mathrm{th}}\in[1,M]}\frac{1}{2}\left[P(n<n_{\mathrm{th}}\mid i=1)+P(n\geq n_{\mathrm{th}}\mid i=0)\right].$ (61)
For comparison, we also evaluated the error probability given by the full MAP decision rule:
$P_{\mathrm{err,MAP}}^{\mathrm{ISK}}=1-\frac{1}{2}\sum_{n=0}^{M}\max\!\left\{P(n\mid i=0),\,P(n\mid i=1)\right\}.$ (62)
We found that these two error probabilities coincide numerically over the parameter ranges considered in this work, indicating that, within the explored regimes, the full MAP rule is effectively equivalent to an optimal threshold rule.
Simulation implementation
All numerical simulations were performed in Python using QuTiP (version: 5.1.1) [45], together with NumPy and SciPy. QuTiP was used to construct the relevant single-mode quantum states and operators in a truncated Fock basis, including displaced squeezed states, coherent states, thermal states, and the corresponding transformed states after displacement, inverse squeezing, and noisy channels.
For the phase-diffusion model, the mixed received states $\rho_{i}^{\text{pd}}$ were generated by averaging over random phase rotations weighted by the Gaussian phase-noise distribution, see Eq. (35). For the thermal-noise model, the states $\zeta_{i}^{\prime}$ entering the PNR(M) were obtained from the OOK displacement (i.e. $D(0)$ and $D(2\gamma)$) representation of the thermal state $\rho_{\text{th}}(n_{t})$, see Eq. (43). For the pure-loss channel, the received states $\rho_{i}^{\text{loss}}$ were modeled as BPSK displaced squeezed thermal states, see Eq. (51). In all noisy cases, the conditional counting statistics were evaluated numerically and inserted into the MAP rule to determine the optimal decision and corresponding error probability.
The photon-number cutoff $N_{\text{cut}}$ in the Fock basis was chosen sufficiently large to ensure numerical convergence for all reported parameter ranges. In this paper, the simulations use a cutoff $N_{\text{cut}}=100$. We verified numerical convergence by increasing the cutoff dimension until the resulting error probabilities changed negligibly. The same convergence criterion was applied to the calculated Wigner functions and Fock-basis populations shown in the figures.
Acknowledgements
@@
Data Availability
The data that support the findings of this study are available from the authors upon request.
Code Availability
The numerical simulations in this study were performed in Python using QuTiP 5.1.1. The custom scripts used to generate the results are available from the corresponding author on reasonable request.
@@
The Authors declare no Competing Financial or Non-Financial Interests.
Figure 1: Comparison of the Helstrom bound (HB) and the standard quantum limit (SQL) for error probabilities in binary phase-shift keying (BPSK) with displaced squeezed states (S-BPSK) and coherent states (C-BPSK). Error probability is plotted against the signal mean photon number $N$. The C-BPSK Helstrom bound $P_{\text{HB}}^{\text{C-BPSK}}=\frac{1}{2}\left[1-\sqrt{1-\exp\left(-4\cdot N\right)}\right]$ is used as the benchmark, and the corresponding standard quantum limit is $P_{\text{SQL}}^{\text{C-BPSK}}=\frac{1}{2}\left[1-\mathrm{erf}\left(\sqrt{2\cdot N}\right)\right]$.
[…]
Figure 3: Evolution of received states in the proposed inverse-squeezing Kennedy receiver. Panels (a)-(c) show phase-space representations, while panels (d)-(f) show the corresponding Fock-basis populations. Panels (a) and (d) depict the received states $\rho_{i}$. Panels (b) and (e) display the states after the nulling displacement $\sigma_{i}=D(\alpha)\rho_{i}D^{\dagger}(\alpha)$. Panels (c) and (f) present the states after inverse-squeezing $\zeta_{i}=U\rho_{i}U^{\dagger}$ with $U=S(-r)D(\alpha)$. The mean photon number is $N=1.0$ for all panels.
Figure 4: Performance of the proposed inverse-squeezing Kennedy receiver (IS-Kennedy) under ideal conditions. (a) Error probability of the IS-Kennedy ($P_{\text{err}}^{\text{ISK}}$). (b) Ratio (in dB) of the error probability limits (Helstrom bound (HB) and standard quantum limit (SQL)) of binary phase-shift keyed displaced squeezed states (S-BPSK) and coherent states (C-BPSK) to $P_{\text{err}}^{\text{ISK}}$ (as the benchmark).
Figure 5: Performance of the proposed inverse-squeezing Kennedy receiver (IS-Kennedy) using an imperfect photon-number-resolution (PNR) detection with finite photon number resolution M (PNR(M)). (a) Error probability of IS-Kennedy with PNR(1) (i.e. single-photon detector) for different detection efficiencies $\eta\in\left\{1.0,0.8,0.4\right\}$ and extremely small dark counts $\nu=10^{-9}$ (thick solid); the error probability of Kennedy receiver with $\eta=1,\ \nu=10^{-9}$ (solid purple); the Helstrom bound (HB) of binary phase-shift keyed (BPSK) coherent states (C-BPSK) (dot-dash blue); and the standard quantum limit (SQL) of BPSK squeezed states (S-BPSK) (short-dash yellow). (b) Error probability of IS-Kennedy with PNR(2) for different dark counts $\nu\in\left\{10^{-2},10^{-3},10^{-5}\right\}$ and ideal detection efficiency $\eta=1$ (thick solid). (c) The optimal detection threshold $n_{\text{th}}^{*}$ as a function of signal energy for different dark counts $\nu\in\left\{10^{-2},10^{-3},10^{-5}\right\}$ (solid). The jumps in threshold correspond exactly to the inflection points observed in panels (b) and (d). (d) The ratio (in dB) of the IS-Kennedy with PNR(10) (solid) to the S-BPSK SQL (as the benchmark, short-dash yellow). The oscillatory behavior reflects the discrete adjustment of $n_{\text{th}}^{*}$ to balance signal detection against dark count noise.
Figure 6: Schematic of the proposed inverse-squeezing Kennedy receiver (IS-Kennedy) under phase diffusion channel. The signal state $\rho_{i}$ (pure state) is transformed into the state $\rho_{i}^{\text{pd}}$ (mixed state) in the phase diffusion channel.
Figure 7: The evolution of phase-space representations of binary phase-shift keyed (BPSK) displaced squeezed states (S-BPSK). (a) S-BPSK $\rho_{i}$ without phase diffusion noise; (b) states $\rho_{i}^{\text{pd}}$ with phase diffusion noise; (c) states $\zeta_{i}$ after the processing of the proposed receiver. Parameters for the panels: mean photon number $N=1.0$ and phase diffusion strength $\sigma=0.5$.
Figure 8: Simulation results of the proposed inverse-squeezed Kennedy (IS-Kennedy) using photon-number-resolving (PNR) detection with finite photon number resolution M (PNR(M)) under phase diffusion noise. (a) Error probability of IS-Kennedy for different resolution $\text{M}\in\{1,2,3,4,10\}$ (thick solid), the Helstrom bound (HB, short-dash brown) and standard quantum limit (SQL, short-dot blue) of binary phase-shift keyed (BPSK) displaced squeezed states(S-BPSK). (b) Ratio (in dB) of IS-Kennedy (thick solid) to the S-BPSK SQL (as the benchmark, short-dot blue) for different phase diffusion strength $\sigma\in\{0.01,0.1,0.2,1.0\}$, (c) Optimal threshold $n_{th}^{*}$ for a resolution of 10, and (d) Fock-basis population of measured states $\zeta_{i}=U\rho_{i}^{\text{pd}}U^{\dagger}$ for $N=1.5$. The analytical expressions for the SQL and HB of S-BPSK under phase diffusion are provided in Supplementary Information II. Parameters for the panels (a) (c) and (d): phase diffusion strength $\sigma=0.5$.
Figure 9: Schematic of the proposed inverse-squeezing Kennedy receiver (IS-Kennedy) under receiver thermal noise. The measured states $\zeta_{i}$ impinging on the photon-number-resolving (PNR) detection with finite photon number resolution M (PNR(M)) are contaminated by receiver thermal noise and become $\zeta_{i}^{\prime}$.
Figure 10: Simulation results of the proposed inverse-squeezing Kennedy receiver (IS-Kennedy) using the photon-number-resolving (PNR) detection with finite photon number resolution M (PNR(M)) under thermal noise condition. (a) Error probability of IS-Kennedy for different resolutions $\text{M}\in\{1,2,3\}$ (noise strength $n_{t}=0.001$) and the error probability limits (Helstrom bound (HB, short-dash blue) and standard quantum limit (SQL, dot-dash pink)) of binary phase-shift keyed (BPSK) displaced squeezed states (S-BPSK). (b) Error probability of the IS-Kennedy ($\text{M}=10$, thick solid) for $n_{t}\in\{10^{-3},10^{-2},0.1,0.2,1.0\}$. (c) Ratio (in dB) of the IS-Kennedy ($\text{M}=10$, thick solid) to the S-BPSK SQL (as the benchmark, dot-dash pink) for $n_{t}\in\{10^{-3},10^{-2},0.1,0.2,1.0\}$. (d) Optimal threshold of the IS-Kennedy ($\text{M}=10$, thick solid) for $n_{t}\in\{10^{-3},10^{-2},0.1,0.2,1.0\}$. (e) Fock-basis population of $\zeta_{i}^{\prime}$ (the input energy $N=1.0$ and $n_{t}=0.001$), and (f) Maximum tolerable thermal noise $n_{t}^{\text{max}}$ for $\text{M}\in\{1,2,3,4\}$.
Figure 11: Schematic of the proposed inverse-squeezing Kennedy receiver (IS-Kennedy) with the pure-loss channel. The signal state $\rho_{i}$ (pure state) is transformed into the state $\rho_{i}^{\text{loss}}$ (mixed state) in the pure-loss communication channel, which $T$ is the equivalent transmittance.
[…]